OpenStudy (anonymous):

solve for w in P=2w + 2l, if P=38 and l=12

OpenStudy (anonymous):

you should take p and I into their places

OpenStudy (anonymous):

substitute value of 'p' and the value of 'l'... 38 = 2w + 2 (12) now solve for w...

OpenStudy (texaschic101):

P = 2w + 2l (when p = 38 and l = 12) sub 38 in for p and 12 in for l P = 2w + 2l 38 = 2w + 2(12) can you work it out from here ?

OpenStudy (anonymous):

it's a simple work you need to subtract 2(12) from both side of the equation

OpenStudy (anonymous):

it will be like this 38 - 2(12) = 2w + 2(12) - 2(12)

OpenStudy (anonymous):

W=7